OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) − 4

sam (.sam.):

\[\begin{array}{l} 3 (x+1)=-2 (x-1)-4 \\ \text{expand out terms on both sides} \\ 3 x+3=-2 x-2 \\ \text{add }2 x-3\text{ to both sides} \\ 5 x=-5 \\ \text{the last step is to divide both sides by 5} \\\end{array}\]